### v1

# Limitations of Using Identical Distributions for Training and Testing When Learning Boolean Functions

Abstract
When the distributions of the training and test data do not coincide, the problem of understanding generalization becomes considerably more complex, prompting a variety of questions. In this work, we focus on a fundamental one: Is it always optimal for the training distribution to be identical to the test distribution? Surprisingly, assuming the existence of one-way functions, we find that the answer is no. That is, matching distributions is not always the best scenario, which contrasts with the behavior of most learning methods. Nonetheless, we also show that when certain regularities are imposed on the target functions, the standard conclusion is recovered in the case of the uniform distribution.
1 Introduction
When analyzing theoretically supervised learning methods, assumptions about the problem are usually made. Typically, we assume that the training distribution coincides with the test distribution, or that there exists some true labeling rule for the given data. However, in real scenarios, some of these assumptions are not satisfied, raising the question of how this affects the performance of the learning methodologies.
Deviations of the test distribution from the training distribution usually degrade the performance of the methods used [hand2006classifier, daume2006domain]. For this reason, alternative methods that provide some protection against such distributional shifts have been proposed. For instance, [duchi2021learning, blanchet2019robust] study the use of a worst-case type of loss, where instead of minimizing the empirical risk, one minimizes the supremum of the risk over a set of distributions close to the training distribution. This approach ensures that small distributional shifts do not produce significant changes in the performance of the learning method. Other approaches, such as [montasser2024transformation], consider different types of shifts in the test distribution, where the test data has the form $(T(x),y)$, with $(x,y)$ distributed as in the training set, and $T$ being some function. The intuition here is that, for instance, when applying rotations to images in a classification task, the label should remain unchanged. Domain adaptation methods [ganin2015unsupervised, fernando2013unsupervised, shimodaira2000improving, ben2006analysis, ben2010theory], used mainly for covariate shift, i.e., when the labeling rule remains unchanged but only the marginal distribution of the input variable changes, make use of (usually unlabeled) data from the test distribution to adapt the model to the new distribution.
In this work, however, we do not attempt to establish under which shifts certain learning methods are robust, but rather address a more fundamental question: is it really optimal for the test distribution to be identical to the training distribution? To investigate this, we use tools from computational learning theory. In particular, instead of adopting the traditional PAC-learning framework introduced by Valiant [valiant1984theory], we focus on a variant in which the efficient evaluation of the model is also considered [gyurik2023exponential, perez2024classical]. Intuitively, one might think the answer to the question is trivially “yes.” However, we show that this is not generally the case, assuming that one-way functions exist, which highlights the extremely complex and counterintuitive nature of understanding generalization in this setting. Nonetheless, assuming certain regularities, we can guarantee that this effect disappears for the uniform distribution.
This paper is organized as follows. Section 2 introduces the necessary notation and definitions used throughout the paper. Next, in Section 3, the main results are presented, showing that even if a concept class can be learned under some distribution, it may not be learnable when the training distribution coincides with the test distribution. In Section 4, it is shown that, at least for the uniform distribution, if the concept class satisfies certain regularity conditions, the aforementioned disparities disappear. Finally, in Section 5, the conclusions are presented.
2 Notation and Definitions
In this paper, we analyze the scenario of learning Boolean functions using a variant of the PAC-learning framework. To proceed, we introduce the necessary notation and definitions. First, in this setting, we are given the side information that the target function belongs to some concept class $\mathcal{C}_{n}$. Therefore, we aim to design a learning algorithm $\mathcal{A}$ that estimates any Boolean function $c_{n}:\{0,1\}^{n}\rightarrow\{0,1\}$ belonging to $\mathcal{C}_{n}$, given a training set $\mathcal{T}_{S}^{c_{n}}=\{(x_{i},c_{n}(x_{i}))\}_{i=1}^{S}$, where $x_{i}\sim D^{T}_{n}$. Specifically, algorithm $\mathcal{A}$ has two inputs: $x$ and a training set $\mathcal{T}_{S}^{c_{n}}$, and outputs an estimate of $c_{n}(x)$. That is, it can be understood as the concatenation of, first, a learning method that selects a hypothesis $h_{n}:\{0,1\}^{n}\to\{0,1\}$, and then a method that evaluates it on the input $x$. The performance of algorithm $\mathcal{A}$ is measured by the risk, given by
$R(\mathcal{A},c_{n}):=\mathbb{E}_{x\sim D^{E}_{n},\,\mathcal{T}_{S}^{\,c_{n}},\,\mathcal{A}}\left[\ell(\mathcal{A}(x,\mathcal{T}_{S}^{\,c_{n}}),c_{n}(x))\right]$ (1)
where $\ell:\mathcal{Y}_{n}\times\mathcal{Y}_{n}\rightarrow\mathbb{R}^{+}$ denotes the loss function, and $D_{n}^{E}$ is the test distribution. Throughout this paper, we consider $\ell(y,y^{\prime}):=\mathbbm{1}\{y\neq y^{\prime}\}$. In this case, the risk coincides with the error probability. Importantly, when an algorithm appears as a subscript of an expectation, as in (1), this indicates that the expectation is also taken with respect to the algorithm’s internal randomness.
In this setting, the learning problem is analyzed for a sequence of concept classes $\mathcal{C}=\{\mathcal{C}_{n}\}_{n\in\mathbb{N}}$, with an associated sequence of training distributions $D^{T}:=\{D^{T}_{n}\}_{n\in\mathbb{N}}$ and test distributions $D^{E}:=\{D^{E}_{n}\}_{n\in\mathbb{N}}$. With a slight abuse of terminology, we refer to both $\mathcal{C}_{n}$ and $\mathcal{C}$ as a concept class, and similarly, on some occasions, to a sequence of distributions $\{D_{n}\}_{n\in\mathbb{N}}$ as a distribution. With these preliminaries established, we can proceed to the formal definition of a learnable concept class $\mathcal{C}$.
Definition 1.
A concept class $\mathcal{C}$ is learnable for a sequence of test distributions $D^{E}$ if there exist an efficient randomized algorithm $\mathcal{A}$, a sequence of training distributions $D^{T}$, and a multivariate polynomial $p(a,b)$ such that algorithm $\mathcal{A}$, given input $x\in\{0,1\}^{n}$, a Boolean string of $m$ ones $1^{m}$, and a training set $\mathcal{T}_{p(n,m)}^{\,c_{n}}=\{(x_{i},c_{n}(x_{i}))\}_{i=1}^{p(n,m)}$ with $x_{i}\sim D_{n}^{T}$, outputs $\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{\,c_{n}},1^{m})\in\{0,1\}$ satisfying
$\mathbb{E}_{x\sim D_{n}^{E},\,\mathcal{T}_{p(n,m)}^{\,c_{n}},\textnormal{ }\mathcal{A}}\left[\mathbbm{1}\left\{\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{\,c_{n}},1^{m})=c_{n}(x)\right\}\right]\geq 1-\frac{1}{m}$ (2)
for all $n,m\in\mathbb{N}$ and $c_{n}\in\mathcal{C}_{n}$. The expected value is taken with respect to $x\sim D_{n}^{E}$, the training set $\mathcal{T}_{p(n,m)}^{c_{n}}$, and the internal randomness of algorithm $\mathcal{A}$.
In this definition, the algorithm $\mathcal{A}$ takes an additional argument $1^{m}$, which represents a given precision requirement. As is common in computational complexity, we use $1^{m}$ instead of $m$ directly to ensure that any efficient algorithm runs in time $\mathrm{poly}(n,m)$ rather than $\mathrm{poly}(n,\log m)$. We refer to a sequence of training distributions $D^{T}$ as a sufficiently informative distribution for learning $\mathcal{C}$ with respect to $D^{E}$ if there exists an algorithm $\mathcal{A}$ and a polynomial $p(a,b)$ that satisfy the conditions specified in Definition 1. The set of sufficiently informative distributions is denoted by $\mathrsfso{D}(\mathcal{C},D^{E})$.
Finally, we conclude by introducing the definitions of some complexity classes employed in deriving the results of this work: $\mathsf{HeurBPP/samp}(D)$ and $\mathsf{HeurP/poly}$.
Definition 2.
[perez2024classical] $\mathsf{HeurBPP/samp}$(D{D}) is the class of distributional problems $(L,{D}^{\prime})$ that satisfy:
  1. $(i)$
    ${D}^{\prime}$ equals D{D},
  2. $(ii)$
    there exists a probabilistic Turing machine (TM) $M$ running in polynomial time and a polynomial $p(n,m)$, such that for the fixed sequence of distributions ${D}=\{D_{n}\}_{n\in\mathbb{N}}$,
    $\mathbb{P}_{x\sim D_{n}}\left(\mathbb{P}_{\mathcal{T}_{p(n,m)},\,M}\left(M(x,\mathcal{T}_{p(n,m)},1^{m})=\mathbbm{1}\{x\in L\}\right)\geq\frac{2}{3}\right)\geq 1-\frac{1}{m}$ (3)
    for any $n,m\in\mathbb{N}$, where $\mathcal{T}_{p(n,m)}=\{(x_{i},\mathbbm{1}\{x_{i}\in L\})\}_{i=1}^{p(n,m)}$, and $x_{i}\sim D_{n}$. The external probability is taken with respect to $x\sim D_{n}$, and the internal probability with respect to the training set $\mathcal{T}_{p(n,m)}$, and the internal randomness of TM $M$.
Definition 3.
[gyurik2023exponential] $\mathsf{HeurP/poly}$ is the class of distributional problems $(L,D)$ for which there exists a polynomial-time Turing machine $M$ with the following property: for every $n,m\in\mathbb{N}$, there exists a binary string $a_{n,m}\in\{0,1\}^{\mathrm{poly}(n,m)}$ such that
$\mathbb{P}_{x\sim D_{n}}\!\left(\,M(x,a_{n,m},1^{m})=\mathbbm{1}\{x\in L\}\,\right)\geq 1-\frac{1}{m}.$ (4)
Importantly, the condition given in equation (3) can be replaced by
$\mathbb{E}_{x\sim D_{n},\mathcal{T}_{p(n,m)},\textnormal{ }M}\left[\mathbbm{1}\left\{M(x,\mathcal{T}_{p(n,m)},1^{m})=\mathbbm{1}\{x\in L\}\right\}\right]\geq 1-\frac{1}{m}$ (5)
without modifying the definition (see Appendix A of [perez2024classical]). This alternative form will be used throughout the proofs.
3 Limitations of Identical Distributions for Training and Testing
Intuitively, one may think that to learn a concept class for the test distribution $D^{E}$, the best scenario is for the training set to be distributed according to $D^{E}$ as well. However, this intuition is false assuming that one-way functions exist, as shown in the following result.
Theorem 1.
If one-way functions exist, then the following proposition is false: For any pair $(\mathcal{C},D^{E})$ such that there exists a sufficiently informative training distribution $D^{T}$, i.e., $\mathrsfso{D}(\mathcal{C},D^{E})\neq\emptyset$, then $D^{E}\in\mathrsfso{D}(\mathcal{C},D^{E})$.
That is, if one-way functions exist, there are concept classes $\mathcal{C}$ that, although they can be learned for the test distribution $D^{E}$, cannot be learned when the training distribution is the same as the test distribution. The proof follows trivially by combining Lemma 1, Lemma 2, and Lemma 3.
Lemma 1.
If there exists a sequence of distributions $D$ and a language $L$ such that $(L,D)\notin\mathsf{HeurBPP/samp}(D)$ and $(L,D)\in\mathsf{HeurP/poly}$, then the proposition stated in Theorem 1 is false.
Proof.
Let $\mathcal{C}_{n}$ be formed only by the concept $c_{n}(x)=\mathbbm{1}\{x\in L\}$. Then, by the definition of the class $\mathsf{HeurBPP/samp}(D)$, we have $D\notin\mathrsfso{D}(\mathcal{C},D)$. Next, since $(L,D)\in\mathsf{HeurP/poly}$, there exists advice $\{a_{n,m}\}_{n,m\in\mathbb{N}}$ and a TM $M$ that, when combined, allows us to compute $c_{n}(x)$ on average. We now design a distribution that encodes the first $\min\{\ell(a_{n,m}),2^{n}\}$ bits of the advice $a_{n,m}$, where $\ell(z)$ denotes the length of string $z$. Note that although $\ell(a_{n,m})=O(\mathrm{poly}(n,m))$, when $m$ is sufficiently large the length of the string may be greater than $2^{n}$. The bits are encoded as follows: the first bit is encoded in the probability of obtaining the outcome $x=0$. If the first bit is 0, this probability is $p_{0}$, otherwise, it is $2p_{0}$. The second bit is encoded analogously using the probability of the outcome $x=1$, and so on. The value $p_{0}$ is determined by the restriction that the sum of the probabilities is equal to $1$. Finally, as shown in Theorem 1 [perez2024classical], using a polynomial number of samples in $\min\{\ell(a_{n,m}),2^{n}\}$, we can decode all bits correctly with high probability.
If $\ell(a_{n,m})<2^{n}$, then we use the decoded advice $a_{n,m}$ together with the Turing machine $M$ to correctly learn the concept class. By contrast, if $\ell(a_{n,m})>2^{n}$, then with high probability the training set contains all $2^{n}$ possible pairs $(x,c_{n}(x))$. Therefore, to estimate $c_{n}(x)$, we only need to look up its value in the training set. Altogether, this implies that $\mathrsfso{D}(\mathcal{C},D)$ is nonempty, although $D\notin\mathrsfso{D}(\mathcal{C},D)$. ∎
Lemma 2.
If there exists some concept class $\mathcal{C}$ such that:
  1. $(i)$
    At least one sequence of distributions $D$ satisfies $D\notin\mathrsfso{D}(\mathcal{C},D)$.
  2. $(ii)$
    All Boolean functions $c_{n}^{(j)}\in\mathcal{C}_{n}$ can be efficiently computed by an algorithm, i.e., there exists an algorithm that, given $j$ and $x\in\{0,1\}^{n}$, outputs $c_{n}^{(j)}(x)$ in polynomial time.
  3. $(iii)$
    $|\mathcal{C}_{n}|\leq 2^{q(n)}$ for all $n\in\mathbb{N}$, where $q(n)$ is a polynomial.
then, for any sequence of distributions $D$ satisfying $D\notin\mathrsfso{D}(\mathcal{C},D)$, there exists a language $L$ such that $(L,D)\notin\mathsf{HeurBPP/samp}(D)$ and $(L,D)\in\mathsf{HeurP/poly}$.
Proof.
The proof of this lemma is analogous to that of Lemma 1 in [perez2024classical]. For completeness, we provide the proof below. We transform the concept class $\mathcal{C}$ into a language
$L(\mathcal{C}):=\{x:c_{\ell(x)}^{(g(\ell(x)))}(x)=1\}.$ (6)
For the definition of the function $g(n)$, we first introduce some notation. Specifically, given an algorithm $\mathcal{A}$, a sequence of distributions $D$, and a polynomial $p(n,m)$, we define the set $\mathcal{E}_{n}(\mathcal{A},D,p)\subseteq\mathcal{C}_{n}$ as the set of concepts $c_{n}^{(j)}$ that satisfy
$\mathbb{P}_{x\sim D_{n}}\left(\mathbb{P}_{\mathcal{T}_{p(n,m)}^{c_{n}^{(j)}},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{c_{n}^{(j)}},1^{m})=c_{n}^{(j)}(x)\right)\geq\frac{2}{3}\right)<1-\frac{1}{m}$ (7)
for some $m\in\mathbb{N}$, where the samples of the training set are distributed according to $D$. Therefore, by taking $D$ such that $D\notin\mathrsfso{D}(\mathcal{C},D)$, which exists by condition $(i)$, any pair $(\mathcal{A},p)$ satisfies $\mathcal{E}_{n}(\mathcal{A},D,p)\neq\emptyset$ for some $n\in\mathbb{N}$. The set of values of $n\in\mathbb{N}$ for which this occurs is denoted by
${\mathcal{N}}(\mathcal{A},D,p):=\{n\in\mathbb{N}:\mathcal{E}_{n}(\mathcal{A},D,p)\neq\emptyset\}$ (8)
Let $B_{\mathrm{alg}}$ denote a bijection between $\mathbb{N}$ and the set of efficient algorithms, and $B_{\mathrm{poly}}$ represent a bijection between $\mathbb{N}$ and the set of polynomials $\{c\,n^{k_{1}}m^{k_{2}}:(c,k_{1},k_{2})\in\mathbb{N}^{3}\}$. Now, we can define function $g(n)$:
$g(n):=\left\{\begin{matrix}1\text{ if }n\notin\{n_{i}\}_{i=1}^{\infty}\\ \min\{j:c_{n}^{(j)}\in\mathcal{E}_{n}\left(B_{\mathrm{alg}}(s_{i^{*}(n)}),D,B_{\mathrm{poly}}(z_{i^{*}(n)})\right)\}\text{ if }n\in\{n_{i}\}_{i=1}^{\infty}\end{matrix}\right$. (9)
where sequences $s_{i}\in\mathbb{N}$ and $z_{i}\in\mathbb{N}$ satisfy that any point $(a,b)\in\mathbb{N}^{2}$ appears at least once in the sequence $\{(s_{i},z_{i})\}_{i\in\mathbb{N}}$. The sequence $\{n_{i}\}_{i\in\mathbb{N}}$ is defined as follows:
$n_{1}:=\min\{{\mathcal{N}}\left(B_{\mathrm{alg}}(s_{1}),D,B_{\mathrm{poly}}(z_{1})\right)\},$ (10)
and
[…]
Finally, $i^{*}(n)$ denotes the index $i$ such that $n_{i}=n$.
Next, we proceed to prove that $(L(\mathcal{C}),D)\notin\,$$\mathsf{HeurBPP/samp}$($D$), but $(L(\mathcal{C}),D)\in\,\mathsf{HeurP/poly}$. It is straightforward to verify that $(L(\mathcal{C}),D)\in\,\mathsf{HeurP/poly}$. Given $g(n)$, which can be represented with a polynomial in $n$ number of bits (condition $(iii)$), we can efficiently compute $c_{n}^{(g(n))}(x)$ (condition $(ii)$). To establish that $(L(\mathcal{C}),D)\notin\mathsf{HeurBPP/samp}(D)$, we employ a proof by contradiction. That is, we assume that $(L(\mathcal{C}),D)\in\mathsf{HeurBPP/samp}(D)$, which implies the existence of a pair $(\mathcal{A},p)$ such that
$\mathbb{P}_{x\sim D_{n}}\left(\mathbb{P}_{\mathcal{T}_{p(n,m)}^{c_{n}^{(g(n))}},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{c_{n}^{(g(n))}},1^{m})=c_{n}^{(g(n))}(x)\right)\geq\frac{2}{3}\right)\geq 1-\frac{1}{m}$ (12)
for all $n,m\in\mathbb{N}$. Since each point $(a,b)\in\mathbb{N}^{2}$ appears at least once in the sequence $\{(s_{i},z_{i})\}_{i\in\mathbb{N}}$, there exists an index $i$ such that $(s_{i},z_{i})=(B_{\mathrm{alg}}^{-1}(\mathcal{A}),B_{\mathrm{poly}}^{-1}(p))$. Therefore, for $n=n_{i}$, it follows from the definition of $g(n)$ that
$\mathbb{P}_{x\sim D_{n_{i}}}\left(\mathbb{P}_{\mathcal{T}_{p(n_{i},m)}^{c_{n_{i}}^{(g(n_{i}))}},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n_{i},m)}^{c_{n_{i}}^{(g(n_{i}))}},1^{m})=c_{n}^{(g(n_{i}))}(x)\right)\geq\frac{2}{3}\right)<1-\frac{1}{m}$ (13)
for some $m\in\mathbb{N}$. Consequently, the assumption results in a contradiction, showing that $(L(\mathcal{C}),D)\notin\mathsf{HeurBPP/samp}(D)$. ∎
Lemma 3.
If one-way functions exist, then there exists a concept class $\mathcal{C}$ such that:
  1. $(i)$
    The uniform sequence of distributions satisfies $U\notin\mathrsfso{D}(\mathcal{C},U)$.
  2. $(ii)$
    All Boolean functions $c_{n}^{(j)}\in\mathcal{C}_{n}$ can be efficiently computed by an algorithm, i.e., there exists an algorithm that, given $j$ and $x\in\{0,1\}^{n}$, outputs $c_{n}^{(j)}(x)$ in polynomial time.
  3. $(iii)$
    $|\mathcal{C}_{n}|\leq 2^{q(n)}$ for all $n\in\mathbb{N}$, where $q(n)$ is a polynomial.
Proof.
See Appendix A. ∎
Note that the result of Theorem 1 may appear somewhat artificial, since we are allowing the use of non-uniform sequences of distributions for which there clearly does not exist an efficient algorithm to sample from. We now show that even if we restrict the analysis to sequences of distributions for which there exists an efficient algorithm $G$ that, on input $1^{n}$, generates samples from $D_{n}$, the same result still holds. Let $\mathrsfso{D}^{\text{eff}}(\mathcal{C},D^{E})$ denote the set of sequences of distributions that are sufficiently informative for the pair $(\mathcal{C},D^{E})$ and, additionally, can be efficiently sampled. With this notation introduced, we can now state the second result.
Theorem 2.
If one-way functions exist, then the following proposition is false: For any pair $(\mathcal{C},D^{E})$ such that $D^{E}$ is efficiently samplable, and $\mathrsfso{D}^{\mathrm{eff}}(\mathcal{C},D^{E})\neq\emptyset$, then $D^{E}\in\mathrsfso{D}^{\mathrm{eff}}(\mathcal{C},D^{E})$.
Proof.
Under the assumption that one-way functions exist, and following the construction of the concept class in Lemma 3, there exists a concept class $\mathcal{C}$ such that:
  1. $(i)$
    The uniform sequence of distributions satisfies $U\notin\mathrsfso{D}(\mathcal{C},U)$.
  2. $(ii)$
    All Boolean functions $c_{n}^{(j)}\in\mathcal{C}_{n}$ can be efficiently computed by an algorithm, i.e., there exists an algorithm that, given $j$ and $x\in\{0,1\}^{n}$, outputs $c_{n}^{(j)}(x)$ in polynomial time.
  3. $(iii)$
    $|\mathcal{C}_{n}|\leq 2^{n}$ for all $n\in\mathbb{N}$.
  4. $(iv)$
    For any algorithm $\mathcal{A}$ and any polynomials $p(n,m)$ and $r(n)$,
    $\mathbb{P}_{c_{n}^{(j)}\in\mathcal{C}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{c^{(j)}_{n}},1^{r(n)})=c_{n}^{(j)}(x)\right)\leq\frac{3}{4}+\epsilon(n)$ (14)
    where $\epsilon(n)\rightarrow 0$ as $n\rightarrow\infty$.
In this proof, we modify the concept class $\mathcal{C}$ by altering each concept $c_{n}^{(j)}$ only on the inputs $x=1,2,\ldots,n$ as follows:
$\tilde{c}_{n}^{(j)}(u)=[j]_{u}$ (15)
where $j]_{u}$ denotes the $u^{th}$ bit of $j$. Let $\tilde{\mathcal{C}}_{n}$ denote this modified concept class. Clearly, the uniform distribution over $\{1,\ldots,n\}$ is sufficiently informative for the concept class $\tilde{\mathcal{C}}_{n}$ (this follows from condition ($ii$)), and moreover, it is efficiently samplable. That is, $\mathrsfso{D}^{\mathrm{eff}}(\tilde{\mathcal{C}},U)\neq\emptyset.$
Next, let us denote the random variable $V^{(j)}$ as the pair $(X,c_{n}^{(j)}(X))$ and the random variable $\tilde{V}^{(j)}$ as the pair $(X,\tilde{c}_{n}^{(j)}(X))$, where $X\sim U_{n}$. The total variation distance satisfies
$d_{TV}(V^{(j)},\tilde{V}^{(j)})$ $=\frac{1}{2}\sum_{x\in\{0,1\}^{n},\,y\in\{0,1\}}\left|\mathbb{P}_{(x,y)\sim V^{(j)}}((x,y))-\mathbb{P}_{(x,y)\sim\tilde{V}^{(j)}}((x,y))\right|$
$=\frac{1}{2}\sum_{x\in[n],\,y\in\{0,1\}}\left|\mathbb{P}(x)\mathbbm{1}\{y=c_{n}^{(j)}(x)\}-\mathbb{P}(x)\mathbbm{1}\{y=\tilde{c}_{n}^{(j)}(x)\}\right|$
$=\sum_{x\in[n]}\mathbb{P}(x)\mathbbm{1}\{c_{n}^{(j)}(x)\neq\tilde{c}_{n}^{(j)}(x)\}\leq\frac{n}{2^{n}}$ (16)
Consequently, in order to distinguish between the two random variables, an exponential number of samples is required [lehmann1986testing]. Specifically, to distinguish between these distributions with an average probability of error $\epsilon$ (assuming a prior of $1/2$ for each random variable), we need at least $N\geq\frac{-4\epsilon+2}{d_{TV}\bigl(V^{(j)},\tilde{V}^{(j)}\bigr)}$ samples. This implies that for any polynomial $q(n)$, and for sufficiently large $n$,
$\Big|\mathbb{P}_{c_{n}^{(j)}\in\mathcal{C}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{c^{(j)}_{n}},1^{r(n)})=c_{n}^{(j)}(x)\right)$
$\hskip 85.35826pt-\mathbb{P}_{\tilde{c}_{n}^{(j)}\in\tilde{\mathcal{C}}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{\tilde{c}^{(j)}_{n}},1^{r(n)})=\tilde{c}_{n}^{(j)}(x)\right)\Big|\leq\frac{1}{q(n)}$ (17)
as otherwise, we could distinguish between the two random variables using only a polynomial number of samples. Therefore, by combining (3) and (14),
$\mathbb{P}_{\tilde{c}_{n}^{(j)}\in\tilde{\mathcal{C}}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{\tilde{c}_{n}^{(j)}},1^{r(n)})=\tilde{c}_{n}^{(j)}(x)\right)\leq\frac{3}{4}+\epsilon^{\prime}(n)$ (18)
where $\epsilon^{\prime}(n)\rightarrow 0$ as $n\rightarrow\infty$. That is, the uniform sequence of distributions $U$ is not sufficiently informative, i.e., $U\notin\mathrsfso{D}^{\mathrm{eff}}(\tilde{\mathcal{C}},U)$. ∎
4 Regular Concept Classes and Their Implications
Note that the results of Theorem 2 rely on a scenario in which the concept class $\mathcal{C}$ is such that some pairs $(x,c_{n}(x))$ provide significantly more information than others. This raises a natural question: what if all samples are similarly informative? To address this, we introduce the notion of a regular concept class, which aims to capture the idea that each sample is equally informative.
Definition 4.
Let $x^{a}$ denote the vector $x_{1},\dots,x_{a}]^{T}$, and let $N(x^{a})$ denote the number of distinct elements in the vector $x^{a}$. A concept class $\mathcal{C}$ is regular if, for any pair of sequences of distributions $D$ and $D^{\prime}$, and any pair of polynomials $p(n,m)$ and $p^{\prime}(n,m)$, it holds that
$\mathbb{P}_{x^{p(n,m)},z^{p^{\prime}(n,m)}}\bigl(N(x^{p(n,m)})\leq N(z^{p^{\prime}(n,m)})\bigr)\geq 1-\delta(n,m),$ (19)
where $x_{i}\sim D_{n}$ and $z_{i}\sim D^{\prime}_{n}$, implies that for any sequence of distributions $D^{E}$ and any algorithm $\mathcal{A}$, there exists an algorithm $\mathcal{A}^{\prime}$ such that, for any fixed concept $c_{n}^{(j)}\in\mathcal{C}_{n}$ and any $n,m\in\mathbb{N}$,
$\mathbb{E}_{x\sim D_{n}^{E},\,\mathcal{A}}\left[\mathbbm{1}\left\{\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{c^{(j)}},1^{m})=c_{n}^{(j)}(x)\right\}\right]\leq\mathbb{E}_{x\sim D_{n}^{E},\,\mathcal{A}^{\prime}}\left[\mathbbm{1}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{p^{\prime}(n,m)}^{c^{(j)}},1^{m})=c_{n}^{(j)}(x)\right\}\right]$ (20)
holds with probability at least $1-\delta(n,m)$ over the training sets $\mathcal{T}_{p(n,m)}^{c^{(j)}}=\{x_{i},c_{n}^{(j)}(x_{i})\}_{i=1}^{p(n,m)}$ and $\mathcal{T}_{p^{\prime}(n,m)}^{c^{(j)}}=\{z_{i},c_{n}^{(j)}(z_{i})\}_{i=1}^{p^{\prime}(n,m)}$.
Therefore, in a regular concept class, the number of distinct elements serves as a reliable proxy for the amount of information the training set provides about the concept class $c_{n}$. Consequently, all pairs are expected to be similarly informative about $c_{n}$. Now that we have introduced the regular concept classes, we examine whether a result comparable to Theorems 1 and 2 holds in this context. At least for the uniform distribution, it does not. The formal statement appears next.
Theorem 3.
Let $\mathcal{C}$ be a regular concept class, then if $\mathrsfso{D}(\mathcal{C},U)$ is not empty, then the uniform distribution satisfies $U\in\mathrsfso{D}(\mathcal{C},U)$.
Proof.
See Appendix B. ∎
This result shows that once the concept class satisfies certain regularities, more intuitive conclusions follow. However, it does not imply that the result holds for an arbitrary distribution. That is, even for a regular concept class $\mathcal{C}$, there may exist distributions for which $\mathrsfso{D}(\mathcal{C},D)\neq\emptyset$ does not imply $D\in\mathrsfso{D}(\mathcal{C},D)$.
5 Conclusions
In summary, we have shown that, under the assumption that one-way functions exist, having the training distribution equal to the test distribution is not always optimal. Furthermore, we have shown that this effect appears even when efficiently samplable sequences of distributions are considered. However, to achieve this latter result, the constructed example relies on the fact that some pairs $(x,c_{n}(x))$ are much more informative than others. To study the scenario in which this is not the case, we introduce the concept of a regular concept class, in which all pairs are similarly informative. For this type of concept class, we conclude that, at least for the uniform distribution, the counterintuitive results of Theorem 1 and Theorem 2 do not hold. Finally, as future work, it would be interesting to study other types of concept classes that exhibit the same property as regular concept classes but for a wider set of distributions.
Appendix A Proof Lemma 3
To begin the proof, we use Theorem 3 [goldreich1986construct], which can be restated as:
Theorem 4.
If one-way functions exists, then there exists a sequence of sets of functions $\mathcal{F}_{n}=\{f_{n}^{(j)}\}_{j\in\{0,1\}^{n}}$, where $f_{n}^{(j)}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, such that:
  1. $(i)$
    All Boolean functions $f_{n}^{(j)}$ can be efficiently computed by a classical algorithm, i.e., there exists a classical algorithm such that given $j$ and $x\in\{0,1\}^{n}$ outputs $f_{n}^{(j)}(x)$ in polynomial time.
  2. $(ii)$
    The sequence $\mathcal{F}_{n}$ passes all polynomial-time statistical test for functions.
To understand this theorem, we first need to introduce the notion of a polynomial-time statistical test.
Definition 5.
[goldreich1986construct] A polynomial-time statistical test for functions is a probabilistic polynomial-time algorithm $T$ that, given $n$ as input and access to an oracle $O_{f}$ for a function $f_{n}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$, outputs either 0 or 1. Algorithm $T$ can query the oracle $O_{f_{n}}$ only by writing on a special query tape some $x\in\{0,1\}^{n}$ and will read the oracle answer $f_{n}(x)$ on a separate answer-tape. As usual $O_{f_{n}}$ prints its answer in one step.
We say that a set of functions $\mathcal{F}_{n}$ passes the statistical test $T$ if, for any polynomial $q(n)$,
$\left|\mathbb{P}_{f_{n}^{(j)}\in\mathcal{F}_{n},\,T}\left(T(1^{n},O_{f_{n}^{(j)}})=0\right)-\mathbb{P}_{h_{n}\in\mathcal{H}_{n},\,T}\left(T(1^{n},O_{h_{n}})=0\right)\right|\leq\frac{1}{q(n)}$ (21)
for a sufficiently large $n$. The probabilities are taken with respect to selecting a function uniformly at random from the sets $\mathcal{F}_{n}$ or $\mathcal{H}_{n}$, respectively, as well as over the internal randomness of the test $T$. Here, $\mathcal{H}_{n}$ denotes the set of all Boolean functions $h:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$.
Corollary 4.1.
For any efficient algorithm $\mathcal{A}$, polynomials $p(n,m)$ and $r(n)$, and efficiently samplable sequences of distributions $D^{E}$ and $D^{T}$, it holds that for any polynomial $q(n)$,
$\Big|\mathbb{P}_{f_{n}^{(j)}\in\mathcal{F}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{f^{(j)}_{n}},1^{r(n)})=[f_{n}^{(j)}(x)]_{1}\right)$
$\hskip 85.35826pt-\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=[h_{n}(x)]_{1}\right)\Big|\leq\frac{1}{q(n)}$ (22)
for a sufficiently large $n$, where $x\sim D_{n}^{E}$, $z^{p(n,r(n))}$ denotes p(n,r(n)){p(n,r(n))} i.i.d. samples from $D_{n}^{T}$, and $\mathcal{T}_{p(n,r(n))}^{f^{(j)}_{n}}=\{(z_{i},f_{n}^{(j)}(z_{i}))\}_{i=1}^{p(n,r(n))}$. As before, the notation $f_{n}^{(j)}\in\mathcal{F}_{n}$ and $h_{n}\in\mathcal{H}_{n}$ indicates that the probabilities are taken with respect to selecting a function uniformly at random from the sets $\mathcal{F}_{n}$ or $\mathcal{H}_{n}$, respectively, in addition to the other random variables.
Proof.
Let $T$ be the following polynomial-time statistical test:
  1. (1)
    Take $p(n,r(n))$ samples from the distribution $D^{T}_{n}$, and one sample from $D^{E}_{n}$.
  2. (2)
    Using the oracle, obtain the value of the function at those positions.
  3. (3)
    Compute $\mathcal{A}\bigl(x,\mathcal{T}_{p(n,r(n))}^{f_{n}^{(j)}},1^{r(n)}\bigr)\oplus[f_{n}(x)]_{1},$ where $f_{n}(x)]_{1}$ denotes the first bit of the output of the function $f_{n}$, obtained using the oracle, and $\oplus$ denotes the XOR operation.
Note that since the queries are generated using efficiently samplable distributions, for any polynomial $p(n,m)$ and any efficient algorithm $\mathcal{A}$, the test runs in polynomial time. Consequently, for any polynomial $q(n)$
$\Big|\mathbb{P}_{f_{n}^{(j)}\in\mathcal{F}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{f^{(j)}_{n}},1^{r(n)})=[f_{n}^{(j)}(x)]_{1}\right)$
$\hskip 85.35826pt-\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=[h_{n}(x)]_{1}\right)\Big|\leq\frac{1}{q(n)}$ (23)
holds for a sufficiently large $n$. ∎
To proceed, the following auxiliary result is stated.
Lemma 4.
For any algorithm (even those running in exponential time) $W$, which, on input a training set $\mathcal{T}_{p(n)}^{c_{n}}$ consisting of a polynomial number of pairs, outputs some hypothesis $\hat{c}_{n}$ from the concept class $\mathcal{C}_{n}$ (comprising all possible Boolean functions $c_{n}:\{0,1\}^{n}\rightarrow\{0,1\}$), the following holds for $D_{n}^{T}$ equal to the uniform distribution:
$\mathbb{P}_{c_{n}\in\mathcal{C}_{n},x^{p(n)},W}\left(W(\mathcal{T}_{p(n)}^{c_{n}})\in\mathcal{B}^{\mathcal{C},{D}}_{n}(c_{n},\alpha)\right)\leq\epsilon(n,\alpha)$ (24)
where $\epsilon(n,\alpha)\rightarrow 0$ as $n\rightarrow\infty$, for any $\alpha\in[0,1/2)$, and
$\mathcal{B}^{\mathcal{C},D}_{n}(\tilde{c}_{n},\alpha):=\left\{c_{n}\in\mathcal{C}_{n}:\mathbb{P}_{x\sim D_{n}}\left(c_{n}(x)=\tilde{c}_{n}(x)\right)\geq 1-\alpha\right\}$ (25)
Proof.
First,
$\mathbb{P}_{e}(W)$ $:=\mathbb{E}_{c_{n}}\left[\mathbb{E}_{\mathcal{T}_{p(n)}^{c_{n}}}\left[\mathbb{P}_{W}\left(W(\mathcal{T}_{p(n)}^{c_{n}})\notin\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\right)\right]\right]$
$\geq\mathbb{P}_{e}(W_{\text{opt}})$ (26)
where $W_{\mathrm{opt}}$ denotes the algorithm that minimizes the probability of error. As shown in Appendix 3C [Kay98], the optimal algorithm $W_{\text{opt}}$ can be expressed as
$W_{\text{opt}}(\mathcal{T}_{p(n)})$ $=\operatorname*{argmin}_{c_{n}\in\mathcal{C}_{n}}\sum_{\tilde{c}_{n}\in\mathcal{C}_{n}}\mathbbm{1}\{c_{n}\notin\mathcal{B}^{\mathcal{C},D}_{n}(\tilde{c}_{n},\alpha)\}\mathbb{P}(\tilde{c}_{n}|\mathcal{T}_{p(n)})$
$=\operatorname*{argmin}_{c_{n}\in\mathcal{C}_{n}}\sum_{\tilde{c}_{n}\in\mathcal{C}_{n}}\mathbbm{1}\{\tilde{c}_{n}\notin\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\}\mathbb{P}(\tilde{c}_{n}|\mathcal{T}_{p(n)})$ (27)
Therefore,
$\mathbb{P}_{e}(W_{\text{opt}})$ $=\mathbb{E}_{c_{n}}\left[\mathbb{E}_{\mathcal{T}_{p(n)}^{c_{n}}}\left[\mathbbm{1}\left\{W_{\text{opt}}(\mathcal{T}_{p(n)}^{c_{n}})\notin\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\right\}\right]\right]$
$=\mathbb{E}_{c_{n}}\left[\mathbb{E}_{\mathcal{T}_{p(n)}^{c_{n}}}\left[1-\mathbbm{1}\left\{W_{\text{opt}}(\mathcal{T}_{p(n)}^{c_{n}})\in\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\right\}\right]\right]$
$=\sum_{\mathcal{T}_{p(n)},c_{n}}\mathbb{P}(\mathcal{T}_{p(n)},c_{n})\left(1-\mathbbm{1}\left\{W_{\text{opt}}(\mathcal{T}_{p(n)})\in\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\right\}\right)$
$=\sum_{\mathcal{T}_{p(n)}}\mathbb{P}(\mathcal{T}_{p(n)})\sum_{c_{n}}\mathbb{P}(c_{n}|\mathcal{T}_{p(n)})\left(1-\mathbbm{1}\left\{W_{\text{opt}}(\mathcal{T}_{p(n)})\in\mathcal{B}^{\mathcal{C},D}_{n}(c_{n},\alpha)\right\}\right)$
$=\sum_{\mathcal{T}_{p(n)}}\mathbb{P}(\mathcal{T}_{p(n)})\sum_{c_{n}}\mathbb{P}(c_{n}|\mathcal{T}_{p(n)})\left(1-\mathbbm{1}\left\{c_{n}\in\mathcal{B}^{\mathcal{C},D}_{n}(W_{\text{opt}}(\mathcal{T}_{p(n)}),\alpha)\right\}\right)$
$=\sum_{\mathcal{T}_{p(n)}}\mathbb{P}(\mathcal{T}_{p(n)})\left(1-\sum_{c_{n}\in\mathcal{B}^{\mathcal{C},D}_{n}(W_{\text{opt}}(\mathcal{T}_{p(n)}),\alpha)}\mathbb{P}\left(c_{n}|\mathcal{T}_{p(n)}\right)\right)$
$=\sum_{\mathcal{T}_{p(n)}}\mathbb{P}(\mathcal{T}_{p(n)})\left(1-\max_{c^{\prime}_{n}}\sum_{c_{n}\in\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)}\mathbb{P}\left(c_{n}|\mathcal{T}_{p(n)}\right)\right)$ (28)
In this derivation, we use the notation $\mathcal{T}_{p(n)}$ to denote an arbitrary set of the form $\{(x_{i},y_{i})\}_{i=1}^{p(n)}$, where $x_{i}\in\{0,1\}^{n}$ and $y_{i}\in\{0,1\}$. Now, let’s analyze the posterior probability $\mathbb{P}\left(c_{n}|\mathcal{T}_{p(n)}\right)$. To do this, we use the following expression:
$\mathbb{P}(c_{n}|(x,y))$ $=\mathbb{P}((x,y)|c_{n})\frac{\mathbb{P}(c_{n})}{\mathbb{P}((x,y))}$
$=\mathbb{P}(x)\mathbbm{1}\{y=c_{n}(x)\}\frac{\mathbb{P}(c_{n})}{\mathbb{P}((x,y))}$ (29)
Therefore, if the prior over each concept is uniform, the posterior $\mathbb{P}(c_{n}|(x,y))$ is also uniform over the consistent concepts, i.e., those for which $y=c_{n}(x)$. In general,
[…]
where $C(\mathcal{T}_{p(n)})$ denotes the number of concepts that are consistent with the training set $\mathcal{T}_{p(n)}$. Note that
$2^{2^{n}-p(n)}\leq C(\mathcal{T}_{p(n)})\leq|\mathcal{C}_{n}|=2^{2^{n}}$ (31)
The lower bound follows from the fact that at most $p(n)$ different values of $x$ are fixed by the training set, leaving at least $2^{n}-p(n)$ values unfixed. Consequently,
$\mathbb{P}_{e}(W_{\text{opt}})$ $=\sum_{\mathcal{T}_{p(n)}}\mathbb{P}(\mathcal{T}_{p(n)})\left(1-\max_{c^{\prime}_{n}}\sum_{c_{n}\in\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)}\mathbb{P}\left(c_{n}|\mathcal{T}_{p(n)}\right)\right)$
$\geq\sum_{\mathcal{T}_{p(n)}:C(\mathcal{T}_{p(n)})>0}\mathbb{P}(\mathcal{T}_{p(n)})\left(1-\frac{\max_{c^{\prime}_{n}}|\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)|}{C(\mathcal{T}_{p(n)})}\right)$
$=1-\frac{\max_{c^{\prime}_{n}}|\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)|}{C(\mathcal{T}_{p(n)})}$ (32)
where the inequality follows from the fact that some concepts in $\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)$ might be inconsistent with the training set, and that $\mathbb{P}(\mathcal{T}_{p(n)})=0$ whenever $C(\mathcal{T}_{p(n)})=0$. Next, by the definition of $\mathcal{B}^{\mathcal{C},D}$, for the uniform distribution, we have
$|\mathcal{B}^{\mathcal{C},D}_{n}(c^{\prime}_{n},\alpha)|$ $=\sum_{i=0}^{\left\lceil\alpha 2^{n}\right\rceil}\binom{2^{n}}{i}$
$\leq 2^{2^{n}H\left(\frac{\left\lceil\alpha 2^{n}\right\rceil}{2^{n}}\right)}(\left\lceil\alpha 2^{n}\right\rceil+1)$ (33)
where the inequality follows from expression (12.40) in [cover1999elements], i.e.,
$\binom{a}{b}\leq 2^{aH\left(\frac{b}{a}\right)}$ (34)
Therefore,
$\mathbb{P}_{e}(W_{\text{opt}})\geq 1-2^{2^{n}\left(H\left(\frac{\left\lceil\alpha 2^{n}\right\rceil}{2^{n}}\right)-1\right)+p(n)}(\left\lceil\alpha 2^{n}\right\rceil+1)$ (35)
Lemma 5.
Let $D^{T}_{n}$ and $D^{E}_{n}$ be equal to the uniform distribution $U_{n}$. Then, for any algorithm $\mathcal{A}$ and any polynomials $p(n,m)$ and $r(n)$, it holds that, for the concept class $\mathcal{H}_{n}$ of all Boolean functions $h_{n}:\{0,1\}^{n}\rightarrow\{0,1\}$,
$\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)\right)\leq\frac{3}{4}+\epsilon(n)$ (36)
where $\epsilon(n)\rightarrow 0$ as $n\rightarrow\infty$. The probability is taken with respect to selecting a concept uniformly at random from the set $\mathcal{H}_{n}$, the internal randomness of the algorithm $\mathcal{A}$, the random variable $x\sim D_{n}^{E}$, and the random variable $z^{p(n,r(n))}$, which denotes $p(n,r(n))$ i.i.d. samples from $D_{n}^{T}$.
Proof.
The proof proceeds by contradiction. Specifically, we assume that there exists an algorithm $\mathcal{A}$ and polynomials such that
$\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)\right)\geq\frac{3}{4}+\delta$ (37)
where $\delta$ is a strictly positive constant. Define $W(\mathcal{T}_{kp(n,r(n))}^{h_{n}})=\hat{h}_{n}$ as follows: for each $x$, run $\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})$ $k$ times and set $\hat{h}_{n}(x)$ by majority vote. If $\delta=1/4$, then when estimating the concept for all values of $x\in\{0,1\}^{n}$, no errors will occur, and hence
$\mathbb{P}_{h_{n}\in\mathcal{H}_{n},z^{p(n,r(n))},W}\left(W(\mathcal{T}_{kp(n,r(n))}^{h_{n}})\in\mathcal{B}^{\mathcal{H},D}_{n}(h_{n},0)\right)=1$ (38)
which is a contradiction with Lemma 4. Next, to analyze the case $\delta\in(0,1/4)$, we define the set $\mathcal{H}_{\text{bad},n}\subseteq\mathcal{H}_{n}$ as the set of concepts such that
$\mathbb{P}_{x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)\right)<\frac{3}{4}+\frac{\delta}{2}$ (39)
Therefore,
$\frac{3}{4}+\delta$ $\leq\mathbb{P}(h_{n}\in\mathcal{H}_{\text{bad},n})\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)|h_{n}\in\mathcal{H}_{\text{bad},n}\right)$
$+(1-\mathbb{P}(h_{n}\in\mathcal{H}_{\text{bad},n}))\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)|h_{n}\notin\mathcal{H}_{\text{bad},n}\right)$
$<\left(\frac{3}{4}+\frac{\delta}{2}\right)\,\mathbb{P}(h_{n}\in\mathcal{H}_{\text{bad},n})+1-\mathbb{P}(h_{n}\in\mathcal{H}_{\text{bad},n})$ (40)
which implies that
$\mathbb{P}(h_{n}\in\mathcal{H}_{\text{bad},n})<\frac{4\delta-1}{2\delta-1}<1$ (41)
Let $\mathcal{X}_{\text{bad},n}(h)$ be the set
$\left\{x\in\{0,1\}^{n}:\mathbb{P}_{z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)\right)<\frac{1}{2}(1+\delta)\right\}$ (42)
Therefore, for $h\in\mathcal{H}_{\text{good},n}:=\overline{\mathcal{H}_{\text{bad},n}}$,
$\frac{3}{4}+\frac{\delta}{2}$ $\leq\mathbb{P}(x\in\mathcal{X}_{\text{bad},n}(h))\mathbb{P}_{x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)|x\in\mathcal{X}_{\text{bad},n}(h)\right)$
$+(1-\mathbb{P}(x\in\mathcal{X}_{\text{bad},n}(h)))\mathbb{P}_{x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=h_{n}(x)|x\notin\mathcal{X}_{\text{bad},n}(h)\right)$
$<\frac{1}{2}(1+\delta)\,\mathbb{P}(x\in\mathcal{X}_{\text{bad},n}(h))+1-\mathbb{P}(x\in\mathcal{X}_{\text{bad},n}(h))$ (43)
Consequently,
$\mathbb{P}(x\in\mathcal{X}_{\text{bad},n}(h))<\frac{1-2\delta}{2-2\delta}<\frac{1}{2}$ (44)
Finally, for $x\in\mathcal{X}_{\text{good},n}(h)$, the result of the majority vote, denoted by $\mathcal{A}^{\prime}\big(x,\mathcal{T}_{kp(n,r(n))}^{h_{n}}\big)$, satisfies
$\mathbb{P}_{\,z^{kp(n,r(n))},\,\mathcal{A}^{\prime}}\left(\mathcal{A}^{\prime}(x,\mathcal{T}_{kp(n,r(n))}^{h_{n}})=h_{n}(x)\right)\geq 1-\beta^{k}$ (45)
where $0\leq\beta<1$. Therefore, when listing the values $\mathcal{A}^{\prime}(x,\mathcal{T}_{kp(n,r(n))}^{h_{n}})$ for all $x\in\mathcal{X}_{\text{good},n}(h)$, the probability of making at least one error is given by
$\mathbb{P}_{\,z^{kp(n,r(n))},\,\mathcal{A}^{\prime}}$ $\left(\bigcup_{x\in\mathcal{X}_{\text{good},n}(h)}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{kp(n,r(n))}^{h_{n}})\neq h_{n}(x)\right\}\right)\leq\beta^{k}|\mathcal{X}_{\text{good},n}(h)|$
$\leq 2^{n}\beta^{k}$ (46)
Taking $k=n^{2}$, we have that $2^{n}\beta^{k}\rightarrow 0$ as $n\rightarrow\infty$. Hence, with probability
$\mathbb{P}(h_{n}\in\mathcal{H}_{\text{good},n})\mathbb{P}_{\,z^{kp(n,r(n))},\,\mathcal{A}^{\prime}}\left(\bigcap_{x\in\mathcal{X}_{\text{good},n}(h)}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{kp(n,r(n))}^{h_{n}})=h_{n}(x)\right\}\right)$
$\geq\frac{2\delta}{1-2\delta}\left(1-2^{n}\beta^{n^{2}}\right)$ (47)
we correctly list a fraction $\mathbb{P}(x\in\mathcal{X}_{\text{good},n}(h))>1/2$ of all values of $h_{n}(x)$. That is, the output of $W(\mathcal{T}_{kp(n,r(n))}^{h_{n}})$ belongs to $\mathcal{B}^{\mathcal{C},D}_{n}(h_{n},\alpha)$ with $\alpha<1/2$ with probability at least $\frac{2\delta}{1-2\delta}(1-2^{n}\beta^{n^{2}})$, which contradicts Lemma 4. ∎
Finally, we combine these results to prove Lemma 3.
Proof.
Using Corollary 4.1, we have that for $D^{T}$ and $D^{E}$ being equal to the sequence of uniform distributions,
$\Big|\mathbb{P}_{f_{n}^{(j)}\in\mathcal{F}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{f^{(j)}_{n}},1^{r(n)})=[f_{n}^{(j)}(x)]_{1}\right)$
$\hskip 85.35826pt-\mathbb{P}_{h_{n}\in\mathcal{H}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{h_{n}},1^{r(n)})=[h_{n}(x)]_{1}\right)\Big|\leq\frac{1}{q(n)}$ (48)
Note that the second term coincides with the one analyzed in Lemma 5, that is, the concept class consists of all possible Boolean functions $\{0,1\}^{n}\rightarrow\{0,1\}$. Consequently,
$\mathbb{P}_{f_{n}^{(j)}\in\mathcal{F}_{n},x,\,z^{p(n,r(n))},\,\mathcal{A}}\left(\mathcal{A}(x,\mathcal{T}_{p(n,r(n))}^{f^{(j)}_{n}},1^{r(n)})=[f_{n}^{(j)}(x)]_{1}\right)\leq\frac{3}{4}+\epsilon(n)+\frac{1}{q(n)}$ (49)
for any polynomials $p(n,m)$ and $r(n)$, and any efficient algorithm $\mathcal{A}$. Therefore, the concept class $\mathcal{C}_{n}:=\{[f_{n}^{(j)}]_{1}\}_{j\in\{0,1\}^{n}}$ is not average case learnable for the sequence of uniform distributions $U$ when $U$ is the distribution of the training set. ∎
Appendix B Proof Theorem 3
In this appendix, we provide the proof of Theorem 3, which, for completeness, is restated here.
Theorem 3.
Let $\mathcal{C}$ be a regular concept class, then if $\mathrsfso{D}(\mathcal{C},U)$ is not empty, then the uniform distribution satisfies $U\in\mathrsfso{D}(\mathcal{C},U)$.
The theorem is based on three lemmas: Lemma 6, Lemma 7, and Lemma 8, which are given below.
Lemma 6.
For a random walk of the form $W_{i+1}=W_{i}+A$ with $A=\text{Bern}(p)$, $p>0$, the expected number of steps to reach $k\in\mathbb{N}$ is given by $k/p$.
Proof.
Let r.v. $Y$ be distributed according to
$\mathbb{P}_{Y}(y)$ $=\mathbb{P}(W_{y}=k\text{ and }W_{y-1}=k-1)$
$=\binom{y-1}{k-1}p^{k}(1-p)^{y-k}$ (50)
Therefore,
$\mathbb{E}[Y]$ $=\sum_{y=k}^{\infty}y\binom{y-1}{k-1}p^{k}(1-p)^{y-k}$
$=\frac{p^{k}}{(k-1)!}\sum_{y=k}^{\infty}\prod_{j=1}^{k-1}(y-j)(1-p)^{y-k}$
$=\frac{p^{k}}{(k-1)!}\sum_{y=1}^{\infty}(-1)^{k}\frac{d^{k}}{dp^{k}}\left(1-p\right)^{y-1}$
$=\frac{p^{k}}{(k-1)!}(-1)^{k}\frac{d^{k}}{dp^{k}}\frac{1}{p}=\frac{k}{p}$ (51)
Lemma 7.
For any sequence of distributions $D$, and any polynomials $p(n),q(n)\in\mathbb{N}$, there exists a sufficiently large constant $c\in\mathbb{N}$ such that the polynomial $p^{\prime}(n)=2p(n)q(n)+c$ satisfies
$\mathbb{P}_{x^{p(n)},z^{p^{\prime}(n)}}(N(x^{p(n)})\leq N(z^{p^{\prime}(n)}))\geq 1-\frac{1}{q(n)}$ (52)
for any $n\in\mathbb{N}$, where $x_{i}\sim D_{n}$, $z_{i}\sim U_{n}$, and distribution $U_{n}$ denotes the uniform distribution over $\{0,1\}^{n}$.
Proof.
First, note that
$\mathbb{P}_{Z_{i+1}}\left(N(z^{i+1})=N(z^{i})+1|Z^{i}=z^{i}\right)=1-\frac{N(z^{i})}{2^{n}}\geq 1-\frac{p(n)}{2^{n}}$ (53)
for any $i\in[p(n)]$, or as long as $N(z^{i})\leq p(n)$. Let $n_{0}(p)$ be a constant such that, for all $n\geq n_{0}(p)$, it holds that $p(n)<2^{n}$. Therefore, for $n\geq n_{0}(p)$, the random walk $W_{i+1}=W_{i}+A$ with $A\sim\text{Bern}\left(1-\frac{p(n)}{2^{n}}\right)$ grows slower than $N(z^{i+1})$ until reaching the value $p(n)$. Let the random variables $Y$ and $\tilde{Y}$ be defined by $\mathbb{P}(Y=y)=\mathbb{P}(W_{y}=p(n)\text{ and }W_{y-1}=p(n)-1)$ and $\mathbb{P}(\tilde{Y}=y)=\mathbb{P}(N(Z^{y})=p(n)\text{ and }N(Z^{y-1})=p(n)-1)$, respectively. Therefore, from these definitions, it follows that $\mathbb{E}[\tilde{Y}]\leq\mathbb{E}[Y]$. Hence, using Lemma 6,
$\mathbb{E}[\tilde{Y}]\leq\mathbb{E}[Y]=\frac{p(n)}{1-\frac{p(n)}{2^{n}}}:=g(n)$ (54)
Using Markov’s inequality,
$\frac{1}{K}\geq\mathbb{P}(\tilde{Y}\geq K\mathbb{E}[\tilde{Y}])\geq\mathbb{P}\left(\tilde{Y}\geq Kg(n)\right)$ (55)
That is, we have that, with probability at least 11/K1-1/K, it holds that $N(z^{Kg(n)})\geq p(n)$. Finally, taking $K=q(n)$ shows that 2p(n)q(n)2\,p(n)\,q(n) is sufficient for all $n\geq n_{0}(p)$. For $n<n_{0}(p)$, we can guarantee the same by adding a sufficiently large constant $c$.
Lemma 8.
For any bounded functions $f$ and $g$ with image $[0,B]$, and r.v. $X$ and $Y$,
$\mathbb{E}_{X}[f(X)]\leq\mathbb{E}_{Y}[g(Y)]+B\,\mathbb{P}_{X,Y}(f(X)>g(Y))$ (56)
Proof.
For simplicity, we assume that $X$ and $Y$ are discrete random variables.
$\mathbb{E}_{X}[f(X)]$ $=\sum_{x}f(x)\mathbb{P}_{X}(x)$
$=\sum_{x}f(x)\left(\sum_{y}\mathbb{P}_{Y}(y)\mathbbm{1}\{f(x)>g(y)\}+\mathbb{P}_{Y}(y)\mathbbm{1}\{f(x)\leq g(y)\}\right)\mathbb{P}_{X}(x)$
$=\sum_{x,y}f(x)\mathbb{P}_{X}(x)\mathbb{P}_{Y}(y)\mathbbm{1}\{f(x)>g(y)\}+\sum_{x,y}f(x)\mathbb{P}_{X}(x)\mathbb{P}_{Y}(y)\mathbbm{1}\{f(x)\leq g(y)\}$
$\leq B\,\mathbb{P}_{X,Y}(f(X)>g(Y))+\sum_{x,y}g(y)\mathbb{P}_{X}(x)\mathbb{P}_{Y}(y)$
$=B\,\mathbb{P}_{X,Y}(f(X)>g(Y))+\mathbb{E}_{Y}[g(Y)]$ (57)
Now that the different lemmas have been introduced, we can proceed with the proof of Theorem 3.
Proof.
Since $\mathrsfso{D}(\mathcal{C},U)\neq\emptyset$, there exists an algorithm $\mathcal{A}$, a sequence of distributions $D$, and a polynomial $p(n,m)$ satisfying Definition 1. Next, using Lemma 7, we have that the polynomial $p^{\prime}(n,m)=2p(n,m)m+c$ satisfies
$\mathbb{P}_{x^{p(n,m)},z^{p^{\prime}(n,m)}}(N(x^{p(n,m)})\leq N(z^{p^{\prime}(n,m)}))\geq 1-\frac{1}{m}$ (58)
where $z_{i}\sim U_{n}$. Therefore, since the concept class is regular, there exists an algorithm $\mathcal{A}^{\prime}$ such that
$\mathbb{E}_{x\sim U_{n},\,\mathcal{A}}\left[\mathbbm{1}\left\{\mathcal{A}(x,\mathcal{T}_{p(n,m)}^{c_{n}},1^{m})=c_{n}(x)\right\}\right]\leq\mathbb{E}_{x\sim U_{n},\,\mathcal{A}^{\prime}}\left[\mathbbm{1}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{p^{\prime}(n,m)}^{c_{n}},1^{m})=c_{n}(x)\right\}\right]$ (59)
with probability at least 11/m1-1/m over the training sets $\mathcal{T}_{p(n,m)}^{c_{n}}=\{x_{i},c_{n}(x_{i})\}_{i=1}^{p(n,m)}$ and $\mathcal{T}_{p^{\prime}(n,m)}^{c_{n}}=\{z_{i},c_{n}(z_{i})\}_{i=1}^{p^{\prime}(n,m)}$. Now, we can apply Lemma 8 and the fact that the triplet $(\mathcal{A},D,p)$ satisfies Definition 1 to conclude that
$1-\frac{1}{m}\leq\mathbb{E}_{x\sim U_{n},\,\mathcal{T}_{p^{\prime}(n,m)}^{c_{n}},\,\mathcal{A}^{\prime}}\left[\mathbbm{1}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{p^{\prime}(n,m)}^{c_{n}},1^{m})=c_{n}(x)\right\}\right]+\frac{1}{m}$ (60)
That is,
$1-\frac{1}{m}\leq\mathbb{E}_{x\sim U_{n},\,\mathcal{T}_{p^{\prime}(n,2m)}^{c_{n}},\,\mathcal{A}^{\prime}}\left[\mathbbm{1}\left\{\mathcal{A}^{\prime}(x,\mathcal{T}_{p^{\prime}(n,2m)}^{c_{n}},1^{2m})=c_{n}(x)\right\}\right]$ (61)
Consequently, $U\in\mathrsfso{D}(\mathcal{C},U)$. ∎